OpenStudy (snowcrystal):

Help please I will fan and medal Which of the quadratic functions has the narrowest graph?

OpenStudy (snowcrystal):

\[A.) y= -2x^2\] \[B.) y=5x^2\] \[C.) y= -\frac{ 1 }{ 2 } x^2\] \[D.) y=\frac{ 1 }{ 5 } x^2\]

OpenStudy (snowcrystal):

@HelpBlahBlahBlah

OpenStudy (helpblahblahblah):

i dont know sorry

OpenStudy (snowcrystal):

meh it okay

OpenStudy (snowcrystal):

@iPwnBunnies

OpenStudy (ipwnbunnies):

This is just one of those things you can graph and see. The higher the coefficient of the x^2 term, the skinnier the graph looks.

OpenStudy (ipwnbunnies):

Doesn't matter if it's negative or not. That just means the graph is pointing up or down. I'll draw you the graphs.

OpenStudy (ipwnbunnies):